# Cost-Effective, Low Latency Vector Search with Azure Cosmos DB

### v1

@@ 3.2. Adaptations to the algorithm
Given the limited memory available to the indexer and query processor, Algorithms 2 and  1 would be too slow since full precision vectors can not be cached and require random reads in to the SSD.  [^1] It is hard to re-arrange high-dimensional vectors for spatial locality in queries, so each full precision vector read requires reading a random offset into the index. So we modify the search algorithm to traverse the graph using distance between query and the quantized representation of the vectors which can be cached. As observed previously (Subramanya et al., 2019), this does not significantly impact the convergence rate of greedy search. However, we must re-rank a small set of best candidates found in quantized space using distances to full-precision vectors (see Fig. 5). We configure the index so that a query for top-$10$ entries on a graph with degree $32$ and search list size $L=100$ might touch about 3500 quantized vectors, but only about $50$ full precision vectors.
Indexing in quantized space (mostly)
Inserting a vector $p$ first requires querying for it as described in Algorithm 2 to retrieve the set of visited vertices during search. This can be done entirely in quantized space. The next step is to prune the visited vertex set to get the neighbor set of $p$. It can not be done entirely with quantized vectors because highly compressed quantized vectors can help greedy search navigate to neighborhood of $p$, but may not provide enough precision to precisely identify the nearest neighbors. Prune needs the full precision representation of at least some of the vertices in the visited set. Through experiments over multiple datasets, we determined that using full-precision vectors for the 48 closest vectors to $p$ and Product Quantized vectors for the rest does not cause a significant dip in the quality of the index. This is described formally in Algorithm 8.
@@ 3.3. Design of index terms in Cosmos DB
Cosmos DB stores index terms as key-value pairs in Bw-Tree so they can be read via a single index key lookup or a range scan over keys. There are two kind of index terms stored in Bw-Tree: 1) Inverted terms that map each term (path + document specific value) to a set of ’document ids’ that contain them and 2) Forward terms (introduced with Vector Search) that map each term (path + document id) to any arbitrary value.
@@
We designed a new “Forward Term” type that can have an arbitrary value rather than a bitmap (corresponding to posting lists). The ’Adjacency list terms’ use the new forward term format. The term encodes out-neighbors of the graph vertex representing the vector in the document. This is done by TermKey-EncodedValue consisting of Document Id representing the vector and the TermValue consisting of the list of Document IDs of the out-neighbors (see Figure 4). The new value format supports blind incremental updates to the adjacency list to support fast appends. A new corresponding merge callback procedure is added to process the blind updates and consolidate on to an effective value during Bw-Tree consolidations.
Can Quantized Vectors be persisted as Forward Term? We could have modeled Quantized Data with Forward Term as well (which is a more natural representation) and plan to reconcile this in the future.
Extending Term Design for sharded index. By default, we construct one vector index across all the documents in the replica. In some cases, the user might want to query data that matches a shard key. In such a case, it is inefficient to query the entire vector index filtering for the intended shard key. We instead allow the user to declare a “Sharded $\rm DiskANN$” with a vector index policy to create one $\rm DiskANN$ index per value of the declared shard key present in the replica.
To support such indices, we extend the above term design by prefixing TermKey-EncodedValue with a hash of shard key value. This allows us to access both Quantized and Adjacency terms for a given shard and also co-locates the terms for a shard in a continuous key range making it easier to cache for highly active tenants. Also by encoding each logical shard index as just another set of Bw-Tree keys with a different prefix, the decoupling between logical and physical terms helps Cosmos DB store a long tail of tenants on a single replica.
@@ 3.4. Index construction and maintenance
The vectors are inserted, deleted and replaced lazily in the background while other user transactions are going on. This allows maintaining Cosmos DB’s latency SLAs for transactions while also keeping the graph updated dynamically as the users update the vectors.
@@
Cosmos DB charges RUs for processing vectors upfront during the transaction depending on the number of vectors and size of each vector. Customers thus have a predictable requests throughput for their workloads.
@@ 3.5. Query layer
By default, the query engine scans all documents to compute Vector Distances. When either Flat or Q-Flat index is present, the query layer instead uses them.
If quantized terms are used to compute distance to the query, the query engine finds $k^{\prime}$=quantizedVectorListMultiplier×$\times k$ closest vectors to the query in the quantized space, and reranks them to estimate the true top $k$. Re-ranking is done by loading the documents corresponding to $k^{\prime}$ quantized vectors from the store and re-ordering documents based on distance between the query and the full precision vector.
Figure 5. Re-ranking $10$ closest neighbors retrieved in quantized space using full-precision vectors to find the top $k=4$. quantizedVectorListMultiplier=2.5.
If a $\rm DiskANN$ index is present, then the query engine calls the paginated search API to get the required number of documents required by the re-ranking step. The parameter searchListSizeMultiplier controls the quality of search – higher values return more accurate results at higher latency and RU cost. The $L$ parameter sent to paginated search set to $searchListSizeMultiplier*k$.
@@
Filters in the query are evaluated first followed by the ORDER BY. So for the query listed above, the filter c.category is evaluated first resulting in a compressed bitmap representing the set of documents that satisfy the filter. Further query plan depends on the selectivity of the filter — that is, how many of the database points satisfy the filter. If the selectivity is high, e.g., estimated result size $<1000$, the query iterates over the full vectors corresponding to those documents.
Otherwise, a post-filtering approach based on $\rm DiskANN$ pagination is used. The query layer iteratively paginates $\rm DiskANN$ search over quantized vector until it finds at least quantizedVectorListMultiplier×$\times k$ documents that satisfy the predicate (which can be quickly checked using the compressed bitmap).
@@
Cosmos DB backend requests are limited to 5 seconds. If a query does not complete by this time, it is preempted with a continuation token capturing the query state that the client can use to resume the query. While some queries such as regular ORDER BY queries are streaming, pagination in quantized space is not. The partial results in a vector search cannot be serialized to the continuation token as it can be large in size. So, to handle continuation, the partial results are returned to client and the client needs to re-order the backend results across different continuations.
@@
3.6. Filter-aware Search
Given fast access to the bitmap over documents that satisfy the filter, we can modify $\rm DiskANN$ search to make it more likely to find points that satisfy the filter. This is done by Algorithm 7 in Appendix B, which scales down the distances from the query to vectors that satisfy the filter by $\beta<1.0$ in greedy search iteration.
We can compose sharded and filter-aware $\rm DiskANN$ queries. For a multi-tenant collection configured with sharded vector indexing, the queries with filter over the shard key as well additional filters $f$ can limit search to the relevant shard and use the beta-biased greedy search to optimize for filters $f$.
@@ 4. Evaluation
  • MSTuring: 1 billion Bing queries embedded to 100-dimensional floating point using the Turing AGI v5 model (Zhang et al., 2019), with 100,000 queries from the same distribution.
  • YFCC Dataset: 1 million vectors corresponding to a 192-dimensional CLIP embedding applied to (copyleft) photos and videos in Flickr from the year 2004 until early 2014. This also includes metadata such as camera model, country, year, month. The number of documents per year ranges from  30,000 to  144,000.
Runbooks are long-running sequences of insertion, deletion, and queries to simulate various streaming scenarios (et al., [n.d.]). In this work, our first runbook is based on an expiration time model - where each point is inserted with a randomly selected expiration time, at which point it is deleted - as they are a good proxy for common production scenarios. For expiration time model, we use two instances based on the Wikipedia-10M dataset and the MSTuring-1M dataset to benchmark the recall stability of the vector index. Our second runbook is more adversarial and meant to imitate distribution shift. This runbook instance is based on the MSTuring-10M dataset. The dataset is partitioned into 32 clusters, and points are inserted and deleted in clustered order. For each runbook, the same query dataset is used at each query step.
@@
We use the the following parameters unless noted otherwise. The graph degree is 32 to minimize footprint of the index, and a slack factor of 1.3 is used to reduce number of secondary prunes. We use $L=100$ for index construction. The parallelism for mini-batch inserts is set to 8, since replicas might not regularly get more than 3 cores on shared machines. Each physical partition size limit is 50GB and Bw-tree max chain length is set to 3.
@@ 4.1. Query latency and RU charges
Figure 6 shows the query latency and RU charge for an index of 10 million Wiki-Cohere vectors. Note that p50 latency is under 20 ms for the 10 million index with a recall@10 of 90.83%. Increasing $L$-search gives higher recall, but with increased latency and RU charges. The corresponding plots for 1 million and 100,000 vector indices are shown in Figures 16 and 17 in Appendix F.
@@
We also compare the query cost of Cosmos DB with that of the enterprise grade serverless vector indices such as Pinecone, Zilliz, and DataStax in Table 1. We see that for the 10 million Wiki-cohere index, Cosmos DB has about 41×41$\times$41 ×, 15×15$\times$15 ×, and 2×2$\times$2 × lower query cost than Pinecone (enterprise tier), Zilliz (enterprise tier), and DataStax (standard tier) respectively.
In all these experiments, the index is queried in a warm up phase, following by 5000 queries issued one at a time. The Bw-Tree cache is configured to be large enough to cache the quantized and adjacency list index terms.
RU/query \$/1M RU \$/1M queries Storage (\$)
Cosmos DB 60 \$0.25 \$15 \$22.25
Pinecone (Pinecone, 2025) 32 \$24 \$614 \$11.55
@@
Table 1. Query and monthly storage costs of enterprise-grade serverless vector databases for 1M queries over 10M 768D vectors as of May 5th, 2025. Azure Cosmos DB provides 94.64% recall@10.
Figure 7. p50, p95 and p99 query latencies and RU charges for Wiki-Cohere vector indices over 100K, 1 million and 10 million vectors with $L=100$ which provides 94.64% recall@10.
@@ 4.2. Scale out and Query Aggregation
Cosmos DB can scale up to millions of vectors per partition (subject to the 50 GB of data + index size limit), and scale out to billion-scale vector indices. Figure 9 shows the query latency and RU properties of such collections. The MS Turing 1 billion vector collection spans 50 partitions and 100 million vector collection spans 20 partitions. The server latency is measured per partition and client latency is the end to end latency observed in client side. We use a E64is_v3 Azure VM with 64 vCPUs on the same Azure region as the index to aggregate query results with maximum concurrency and minimum latency. Since the query fans out to large partitions concurrently, the client end-to-end latency is sensitive to the worst latency on the server side. The best latency is achieved by using fewer partitions packed with as many vectors as possible. The same applies to RU charges since the query cost grows logarithmically with the number of vectors in a partition, but linearly with the number of partitions.
@@ 4.3. Ingestion
We use the parameters Ingestion $R=64,L_{build}=50$. Bw-Tree chain length is set to 3. We set the cache to be large enough to hold the vector index terms (excluding full precision vectors).
Figure 10 provides a breakdown of the CPU time spent in each insert. Most of the time is spent accessing the quantized vectors, as expected. Since they are mostly cached, this operation is synchronous and involves traversal of cached Bw-Tree terms. The next two dominant components are the time needed to update and read the adjacency lists, which involve traversal of Bw-Tree and potential clean up of terms. The actual distance computations as well as updating candidate sets in $\rm DiskANN$ is small – under 2ms/insert.
Figure 10. Latency without RG per batch, insertion for MSTuring 1M and breakdown of components (quantized vectors, adjacency list, full vectors), with Bw-Tree chain length = 3
We also consider costs for vector ingestion on the WikiCohere-10M dataset, comparing Azure Cosmos DB with enterprise-grade serverless vector databases Pinecone, Zilliz, and DataStax in Table2. Cosmos DB provides lower vector insertion charges, about $33\%$ and $56\%$ less than Pinecone and DataStax (standard tier) respectively. Cosmos DB does have 5.4×5.4$\times$5.4 × higher insertion charges than Zilliz. However, provisioned throughput and autoscale billing models offer at least a 7×7$\times$7 × price discount on RU charges compared to serverless, so these higher costs can be further reduced.
@@
DataStax(DataStax, 2025) \$0.04 7,680 \$370.2
Table 2. Insertion costs of enterprise-grade serverless vector databases for 10 million 768D vectors (Wiki-Cohere-10M) as of May 5th, 2025.

### v2

@@ 3.2. Adaptations to the algorithm
Given the limited memory available to the indexer and query processor, Algorithms 2 and  1 would be too slow since full precision vectors cannot be cached and require random reads in to the SSD.  [^1] It is hard to re-arrange high-dimensional vectors for spatial locality in queries, so each full precision vector read requires reading a random offset into the index. So we modify the search algorithm to traverse the graph using distance between query and the quantized representation of the vectors which can be cached. As observed previously (Subramanya et al., 2019), this does not significantly impact the convergence rate of greedy search. However, we must re-rank a small set of best candidates found in quantized space using distances to full-precision vectors (see Fig. 5). We configure the index so that a query for top-101010 entries on a graph with degree 323232 and search list size $L=100$ might touch about 3500 quantized vectors, but only about 505050 full precision vectors.
Indexing in quantized space
Inserting a vector $p$ first requires querying for it as described in Algorithm 2 to retrieve the set of visited vertices during search. Prior work has demonstrated that this can be done entirely in quantized space. The next step is to prune the visited vertex set to get the neighbor set of $p$. Graph based indices, as described in prior work, use full precision vectors for the prune stage. We found that during prune, computations can also be done on quantized vectors with moderate compression rates without reducing index quality. For example, while indices over OpenAI text-3-large embeddings can be navigated with 128 byte PQ representations, pruning accurately needs 256 bytes. We use this compression level for indexing and search.
@@ 3.3. Design of index terms in Cosmos DB
Cosmos DB stores index terms as key-value pairs in Bw-Tree so they can be read via a single index key lookup or a range scan over keys. There are two kinds of index terms stored in Bw-Tree: 1) Inverted terms that map each term (path + document specific value) to a set of ’document ids’ that contain them and 2) Forward terms (introduced with Vector Search) that map each term (path + document id) to any arbitrary value.
@@
We designed a new “Forward Term” type that can have an arbitrary value rather than a bitmap (corresponding to posting lists). The ’Adjacency list terms’ use the new forward term format. The term encodes out-neighbors of the graph vertex representing the vector in the document. This is done by TermKey-EncodedValue consisting of Document Id representing the vector and the TermValue consisting of the list of Document IDs of the out-neighbors (see Figure 4). The new value format supports blind incremental updates to the adjacency list to support fast appends. A new corresponding merge callback procedure is added to process the blind updates and consolidate on to an effective value during Bw-Tree consolidations. It would have been more natural to model Quantized Data with Forward Term as well, and plan to reconcile this in the future.
Extending Term Design for sharded index. By default, we construct one vector index across all the documents in the replica. In some cases, the user might want to query data that matches a shard key. In such a case, it is inefficient to query the entire vector index filtering for the intended shard key. We instead allow the user to declare a “Sharded $\rm DiskANN$” with a vector index policy to create one $\rm DiskANN$ index per value of the shard key present in the replica.
To support such indices, we extend the above term design by prefixing TermKey-EncodedValue with a hash of shard key value. This allows us to access both Quantized and Adjacency terms for a given shard and also co-locates the terms for a shard in a continuous key range making it easier to cache for highly active tenants. By encoding each logical shard index as just another set of Bw-Tree keys with a different prefix, the decoupling of logical and physical terms helps store a long tail of tenants on a single replica.
@@ 3.4. Index construction and maintenance
Updates to the graph index to reflect insertions, deletions and replications are performed outside the transactional path to allow Cosmos DB to maintain latency SLAs for transactions. We calibrate the charges for transactions with vector operations to rate limit them so that the graph index running in the background can catch up to transactions.
@@
Cosmos DB charges RUs for processing vectors upfront during the transaction depending on the number of vectors and size of each vector to offer predictability to users.
@@ 3.5. Query layer
[…]
Figure 5. Re-ranking 101010 closest neighbors retrieved in quantized space using full-precision vectors to find the top $k=4$. quantizedVectorListMultiplier=2.5.
If a $\rm DiskANN$ index is present, then the query engine calls the Paginated search API to get the number of documents required by the re-ranking step. The parameter searchListSizeMultiplier controls the quality of search – higher values return more accurate results at higher latency and RU cost. The $L$ parameter sent to paginated search set to $searchListSizeMultiplier*k$.
@@
[…]
Filter Aware Graph Search
Given fast access to the bitmap over documents that satisfy the filter, we can modify $\rm DiskANN$ search to make it more likely to find points that satisfy the filter. This is done by Algorithm 7 in Appendix B, which scales down the distances from the query to vectors that satisfy the filter by $\beta<1.0$ in greedy search iteration. With this method, we can support the broad set of predicates that Cosmos DB supports with filter aware graph search.
Highly Selective Queries
Queries with predicates that match fewer than 5000 documents are routed by the query planner to use either the quantized flat index or full precision vectors in the document. Query planned uses the quantized vectors stored in Bw-tree which supports fast range scan when the document matching the filters are in a relatively small range. For queries that have match few documents scatter across large range, query planner decides to load individual points from document store instead of doing a large range quant vector scan.
Sharded DiskANN
For larger indices with suitable distributions on their labels, we require a sharded $\rm DiskANN$ index policy to provide shard key base on label. This allows us to create $\rm DiskANN$ instance base on the shard key and a filtered query to route directly to the shard corresponding to its label, or one of its labels. We can compose sharded and filter-aware $\rm DiskANN$ queries. For a multi-tenant collection configured with sharded vector indexing, the queries with filter over the shard key as well additional filters $f$ can limit search to the relevant shard and use the beta-biased greedy search to optimize for filters $f$.
@@
Cosmos DB backend requests are limited to 5 seconds. If a query does not complete by this time, it is preempted with a continuation token capturing the query state that the client can use to resume the query. While some queries such as regular ORDER BY queries are streaming, pagination in quantized space is not. The partial results in a vector search cannot be serialized to the continuation token as it can be large in size. So, to handle continuation, the partial results are returned to client, which is responsible for reordering and merging them across continuations.
@@ 4. Evaluation
  • MSTuring: 1 billion Bing queries embedded to 100 dimensional floating point using the Turing AGI v5 model (Zhang et al., 2019), with 100,000 queries from the same distribution.
  • YFCC Dataset: 1 million vectors corresponding to a 192 dimensions CLIP embedding applied to (copyleft) photos and videos in Flickr from the year 2004 until early 2014. This also includes metadata such as camera model, country, year, month. The number of documents per year ranges from  30,000 to  144,000.
Runbooks are long-running sequences of insertion, deletion, and queries to simulate various streaming scenarios (Simhadri et al., 2024a, b). In this work, our first runbook is based on an expiration time model - where each point is inserted with a randomly selected expiration time, at which point it is deleted - as they are a good proxy for common production scenarios. For expiration time model, we use two instances based on the Wikipedia-10M dataset and the MSTuring-1M dataset to benchmark the recall stability of the vector index. Our second runbook is more adversarial and meant to imitate distribution shift. This runbook instance is based on the MSTuring-10M dataset. The dataset is partitioned into 32 clusters, and points are inserted and deleted in clustered order. For each runbook, the same query dataset is used at each query step.
@@
We use the the following parameters unless noted otherwise. The graph degree is 32 to minimize footprint of the index, and a slack factor of 1.3 is used to reduce number of secondary prunes. We use $L=100$ for index construction. The parallelism for mini-batch inserts is set to 8, since replicas might not regularly get more than 3 cores on shared machines. Each physical partition size limit is 50GB and Bw-tree max chain length is set to 15. When measuring query performance, the index is queried in a warm up phase, following by 5000 queries issued one at a time. The Bw-Tree cache is configured to be large enough to cache the quantized and adjacency list index terms.
@@ 4.1. Query latency and RU charges
Figure 6 shows the query latency and RU charge for an index of 10 million Wiki-Cohere vectors. Note that with LSearch=50, p50 latency is under 20 ms for the 10 million index with a recall@10 of 91.83%. Increasing $L$-search gives higher recall, but with increased latency and RU charges. The corresponding plots for 1 million and 100,000 vector indices are shown in Figures 16 and 17 in Appendix F.
@@
We compare Azure Cosmos DB vector search with enterprise-grade serverless vector databases – Pinecone(Pinecone, 2025), Zilliz(Zilliz, 2025), and DataStax(DataStax, 2025) – using their publicly available pricing and documentation. For Azure Cosmos DB, we use query costs determined from experimental tests using default settings. Costs for these other services are reflect in their own cost units (Read Units, vCUs, etc.).
In Table 1, we see that for the 10 million Wiki-cohere index, Cosmos DB has nearly 43×43$\times$43 ×, 12×12$\times$12 ×, and 1.75×1.75$\times$1.75 × lower query cost than Pinecone (enterprise tier), Zilliz (enterprise tier), and DataStax (standard tier) respectively.
Table 1. P99 vector search query and monthly storage costs of enterprise-grade serverless vector DBs for 1 million queries over 10 million 768-dimensional vectors as of July 14, 2025. Cosmos DB provides 94.64% recall@10.
RU per \$ per \$ per Storage
Query 1M RUs 1M queries Cost (\$)
Cosmos DB 70 \$0.25 \$17.50 \$22.25
Pinecone (Pinecone, 2025) 32 \$24 \$768 \$11.55
@@
Figure 7. p50, p95 and p99 query latencies and query RU charges for Wiki-Cohere vector indices over 100K, 1 million and 10 million vectors with $L=100$ which provides 94.64% recall@10.
@@
Figure 9. Comparison of the query latency and cost of post-filtering and filter-aware search on the YFCC dataset.
@@
4.2. Filtered queries
Next, we consider the recall, latency, and the cost of two approaches for filtered queries using the YFCC dataset in Figure 9. We compare the post-filtering search – an unmodified graph search which filters the final answer to return only points satisfying the filter – and filter-aware (Algorithm 7, $\beta=0.3$) approaches.
When the number of database points matching the filter is about 5000, the query planner sends queries invokes the $\rm DiskANN$ index with either post-filtering or filter-aware search. Otherwise, query planner falls back to simpler look ups as described in Section3.
We find that both approaches are capable of providing high recall at low latency and RU cost. Post-filtering search produces slightly higher recall than filter-aware search for the same $L$ value, but it suffers from diminishing returns in recall and high p95 and p99 latency. On the other hand, filter-aware search’s recall can be improved by increasing the search depth ($L$), while still retaining predictable tail latency and RU cost.
For instance, $L=200$ with filter-aware search yields the similar recall as $L=200$ with post-filtering search, but with 10x better p99 latency and 5x less p99 cost. This suggests that filter-aware is better for controlling resource utilization and tail latency.
@@ 4.3. Scale out and Query Aggregation
Cosmos DB can scale up to millions of vectors per partition (subject to the 50 GB limit on data and index), and scale out to billion-scale vector indices. Figure 10 shows the query latency and RU properties of such collections. The MS Turing 1 billion vector collection spans 50 partitions, while the 100 million vector collection spans 20. The server latency is measured per partition and client latency is the end to end latency observed in client side. We use a E64is_v3 Azure VM with 64 vCPUs on the same Azure region as the index to aggregate query results with maximum concurrency and minimum latency. Since the query fans out to large partitions concurrently, the client end-to-end latency is sensitive to the worst latency on the server side. The best latency is achieved by using fewer partitions packed with as many vectors as possible. The same applies to RU charges since the query cost grows logarithmically with the number of vectors in a partition, but linearly with the number of partitions.
@@ 4.4. Ingestion
We now study the vector ingestion component that updates the Bw-tree with index terms reflecting new vector insertions. The parallelism for batch inserts is set to 32 threads, The underlying machine has 16 physical and 32 vCPUs. We set the cache to be large enough to hold the vector index terms in memory (excluding full precision vectors). We set the graph degree $R=32$, index build parameter $L_{build}=100$ and Bw-Tree chain length set to 15. Each insert typically accesses approximately $R\cdot L_{build}$ quantized vectors and just over $L_{build}$ adjacency lists (with a wide margin of error).
Figure 11 provides a breakdown of the CPU time spent in each insert. Most of the time is spent accessing the quantized vectors, as expected. Since they are cached, this operation is synchronous and involves traversal of cached Bw-Tree terms. The next two dominant components are the time spent in DiskANN library and reading the adjacency lists. The actual distance computations as well as updating candidate sets in $\rm DiskANN$ is small – under 3ms/insert.
Figure 11. Latency breakdown for single threaded MS Turing-1M ingestion with components (Quantized vectors, Adjacency list, DiskANN library), Bw-Tree chain length = 15.
Figure 12. Ingestion progress on MS Turing-10M and WikiCohere-10M datasets when Bw-Tree cache is large enough for vector index terms.
[…]
We also evaluate vector ingestion costs on the WikiCohere-10M dataset, using experimental data for Azure Cosmos DB and compare it with the published pricing and cost guidance for the enterprise-grade serverless vector databases their respective costing units equivalent to RUs (i.e., Read Units for Pinecone (Pinecone, 2025) and DataStax (DataStax, 2025) and vCUs for Zilliz (Zilliz, 2025)) in Table 2. Cosmos DB demonstrates lower vector insertion costs—approximately 33% and 53% less than Pinecone and DataStax (standard tier), respectively.
Azure Cosmos DB does have 5.4×5.4$\times$5.4 × higher insertion charges than Zilliz. However, provisioned throughput and autoscale billing models offer at least a 7×7$\times$7 × price discount on RU charges compared to serverless, so these higher costs can be further reduced.
Table 2. Insertion costs of enterprise-grade serverless vector databases for 10 million 768D vectors (Wiki-Cohere-10M) as of July 14th, 2025.
@@
DataStax(DataStax, 2025) \$0.04 768 \$307.2